OpenStudy (anonymous):

Is (dy/dx)^2 equal to (d^2y)/(dx^2)?

OpenStudy (amistre64):

not that I can tell

OpenStudy (bahrom7893):

NO..

OpenStudy (amistre64):

one says to square the derivative; the other says to take the derivative of the derivative

OpenStudy (bahrom7893):

Suppose: dy/dx = x^2+x (dy/dx)^2 = (x^2+x)^2 d^2y/dx^2 = 2x+1

OpenStudy (amistre64):

in some odd case it might be true .... but as a general rule.. no

OpenStudy (bahrom7893):

well yea.. it will be true when: (y')^2 = y'' solve the diff eqn

OpenStudy (anonymous):

I'm a bit of a noob here, I have a follow up question, should I post it here or as a new question?

OpenStudy (amistre64):

new question is always best

OpenStudy (bahrom7893):

new yeah..

OpenStudy (anonymous):

Thanks

OpenStudy (bahrom7893):

cuz we can get EXTRA MEDALS!

OpenStudy (bahrom7893):

haha amistre knows what i mean..

OpenStudy (amistre64):

that and after awhile the post starts to stink; better to keep it fresh :)

OpenStudy (bahrom7893):

it does..